# AI-Generated Images Introduce Invisible Relevance Bias to Text-Image Retrieval

### v3

@@ 3.4. More Serious Bias Caused by Training
Due to wide distribution of AI-generated images on the Internet and the bias in Section 3.3, AI-generated images are very likely to be mixed into the training data of retrieval models. This section delves deeper into the impact on retrieval performance and invisible relevance bias when AI-generated images are mixed into the training of retrieval models. The experimental results reveal a vicious cycle of falling into invisible relevance bias. Specifically, the invisible relevance bias of text-image retrieval models causes the AI-generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of retrieval models, leading to more serious bias and forming a vicious cycle. This ultimately results in users’ search results being surrounded by AI-generated images.
Specifically, we explore the impact of the training mixed with generated images on retrieval by incorporating varying ratios of generated images into the training set. To ensure an accurate assessment, our experiments focus on the model trained from scratch (VSE). This is because pre-trained vision-language models have been pre-trained on massive real text-image pairs, which introduce biases in evaluating the impact of different ratios of generated images in the training data. We reconstruct the training set of Flicker30k by replacing a certain ratio of real images with AI-generated images. For each real image ($I_{r}$) in the training set, we use the method in Section 2 to generate its corresponding AI-generated image ($I_{g}$). Then, we introduce the ratio $\alpha$, which means that in our reconstructed training data, the paired images for $\alpha$ percentage of captions are AI-generated images, and for $(100-\alpha)$ percentage of captions are real images. We change the ratio while keeping the total number of training samples unchanged. We evaluate the performance of the trained model on the test set of Flicker30k+AI (in-domain) and MSCOCO+AI (out-of-domain).
@@
Figure 2. Assessment results on the training set miexed with AI-generated images. We change the ratio of AI-generated images in the datasets (X-axis) while keeping the total number of training samples unchanged. The model is tested on the test set of Flicker30k+AI (in-domain) and MSCOCO+AI (out-of-domain) respectively that we constructed in Section 2.
@@
Figure 2 illustrates that as the ratio increases, the ranking disparity between generated images and real images widens, with the retrieval model exhibiting a greater inclination to rank generated images higher. Meanwhile, Figure 3 demonstrates that with an increasing ratio, the discrepancy in score distribution between generated images and real images increases, and the scores of AI-generated images gradually become greater and greater than those of real images. In both in-domain and out-of-domain settings, the following conclusions can be drawn from these results: (1) Retrieval models trained on the datasets mixed with AI-generated images exhibit more serious invisible relevance bias. (2) The invisible relevance bias tends to become more obvious as the ratio of AI-generated images in the training datasets increases. (3) As the ratio of AI-generated images in the training datasets increases, the retrieval performance on real images is gradually damaged.
@@ 4. Our Method in debiasing
In this section, we propose an effective method to alleviate the invisible relevance bias. The details will be introduced in the following.
@@ 4.1. Design of Debiasing Method
Our method alleviates the retrieval model’s preference bias for generated images by measuring and optimizing the additionally introduced invisible relevance in the training. Given a caption-image pair $(C_{i},I^{r}_{i})$ in the training set, the training process for both dual-encoder and fusion-encoder text-image retrieval models can be formulated as estimating the relevance score $s$ between $C_{i}$ and $I^{r}_{i}$, and using contrastive loss or regression loss as the optimization objective to adjust $s$, which can be described as:
@@
in which $\mathbb{A}$ is the set of all samples in the training data.
@@
(a) $\beta=60\%$ on Flicker30k
(b) $\beta=80\%$ on Flicker30k
(c) $\beta=100\%$ on Flicker30k
(d) $\beta=60\%$ on MSCOCO
(e) $\beta=80\%$ on MSCOCO
(f) $\beta=100\%$ on MSCOCO
Figure 5. T-SNE visualization of the embeddings of real and AI-generated images obtained from the image encoder in VSE.
4.2. Evaluation of Debiasing
We evaluate the effect of our debiasing method from three perspectives: (1) How our method affects the ranking difference between real images and AI-generated images. (2) How our method affects the distribution of caption-image relevance scores. (3) How our method affects the representations of real and AI-generated images. All of these three evaluations are performed in both in-domain (Flicker30k) and out-of-domain (MSCOCO) setting.
Table 5 shows how the $\rm{Relative\triangle}$ and retrieval performance changes with the sampling probability $\beta$. The results indicate that: (1) Our method not only effectively alleviates the retrieval model’s preference for AI-generated images, but also makes real images ranked significantly higher than AI-generated images. (2) As the sampling probability $\beta$ increases, real images are ranked higher and higher than AI-generated images. (3) When $\beta$ is $0.5$, retrieval models can achieve a fair ranking between real and generated images with little bias. (4) Our method improves the retrieval performance. It is because in our method, AI-generated images can be seen as the hard negatives in training, which enhances the ability of the retrieval model to distinguish the images with very similar semantics.
Figure 4 shows the distribution of the caption-image relevance scores estimated by retrieval models with different sampling probability $\beta$. The results indicate that our method effectively reduces the relevance scores between captions and generated images estimated by the retrieval model. With the increase in sampling probability ($\beta$), the disparity in score distribution between the generated images and the real images expands, and the scores of real images gradually become greater and greater than those of generated images.
Figure 5 shows the T-SNE of the embeddings of real and AI-generated images encoded by image encoder in VSE. The reveals that with the increase in the sampling probability ($\beta$), the representations of generated and real images become more distinguishable.
@@ 5. Causes of Invisible Relevance Bias
In this section, we use the debiased model proposed in Section 4 to reversely analyze the causes of the invisible relevance bias. Specifically, we discern the transformations performed by the debiased retriever on AI-generated images through a comparative analysis of the difference in image representations between the original retriever and the debiased retriever. The reverse process of these transformations can be seen as the cause of invisible relevance bias. We also design the experiments to support this point of view.
@@ 5.1. Transformations in Debiasing
To simplify the analysis process, we focus on the dual-encoder retrieval model that estimates the relevance score by computing vector similarity between text and image representations. We use image representation as the main object in our analysis because for the same text, image representation can directly affect the estimation of the relevance score, which can help us to find the most direct causes of invisible relevance bias. Specifically, we analyze the difference between image representations encoded by the original retriever and debiased retriever to get the transformations of debiasing from this difference and reverse these transformations to explore the causes of the invisible relevance bias.
Given the image encoder $v(\cdot;\theta)$ of original text-image retrieval model and image encoder $v^{d}(\cdot;\theta^{\prime})$ of our proposed debiased retrieval model, a AI-generated image $I^{g}$, we can get the representations of $I^{g}$ encoded by $v(\cdot;\theta)$ and $v^{d}(\cdot;\theta^{\prime})$ respectively as:
$\bm{r}=v(I^{g};\theta),\quad\bm{r^{d}}=v^{d}(I^{g};\theta^{\prime}).$
For the normalized text representation $\bm{t}=[t_{1},t_{2},...,t_{n}]$, image representation $\bm{r}=[r_{1},r_{2},...,r_{n}]$ from the original retriever and image representation $\bm{r^{d}}=[r^{d}_{1},r^{d}_{2},...,r^{d}_{n}]$ from the debiased retriever, the relevance $score(\bm{t},\bm{r})$ and $score(\bm{t},\bm{r^{d}})$ can be calculated as:
@@
$\bm{p}$ $=[\triangle r_{1},\triangle r_{2},...,\triangle r_{n}]$
$=[p_{1},p_{2},...,p_{n}],$
@@ 5.2. Reversing the Transformations
We can reverse this process to get the causes of the invisible relevance bias. This is because the transformation made by a debiased retriever to make a biased AI-generated image become unbiased is exactly the difference between the biased AI-generated images and the real images. That is, the reason why an unbiased representation $\bm{r^{d}}$ becomes the representation $\bm{r}$ with bias is that the reverse transformation ($-p_{i}$) is done on $r^{d}_{i}$, which can be described as:
$r_{i}=r^{d}_{i}-p_{i}.$
Therefore, we conclude that an unbiased representaion $\bm{r^{d}}$ becomes the representaion $\bm{r}$ with bias because $\bm{r^{d}}$ is added bitwisely by a vector $\bm{-p}=[-p_{1},-p_{2},-p_{3},...,-p_{n}]$. Combining the phenomenon shown in Figure 6 that there is consistency in the transformations vector $\bm{p}$ on AI-generated images with very different semantics, we can get the causes of invisible relevance bias in AI-generated images: AI-generated images cause the image encoder in the retriever to embed additional information to their representations. This additional information is the direct cause of invisible relevance bias and can be obtained by the difference in image representations between original and debiased retriever. This information has the following three characteristics: (1) This information cannot be reflected in a visible way, but can only be embedded by neural network-based models. (2) When this information is embedded into the representation of the image, it can amplify the query-image relevance to produce a higher score. (3) This information has a certain consistency in AI-generated images with different semantics. This information may be like the watermark that is universal information for the image generation model and can be expressed by neural network visual models such as image encoder. We design the experiments to support these three points.
@@ 6.1. Artificial Intelligence Generated Content
With the development of generation models, AI-generated content (AIGC) becomes more realistic and less discernible (Agnese et al., 2020; Cao et al., 2023; Wu et al., 2023). The use of AI to generate content has greatly changed the way of content generation. It reduced the cost of content generation and improved the efficiency than traditional compared to traditional content generation pipeline biased on humans (Hanley and Durumeric, 2023; Spitale et al., 2023). For example, Stable Diffusion (Yang et al., 2023) and DALL-E-3 (Betker et al., 2023) can generate high-quality images by simply following a short text description of uses’ input. However, as AIGC is gradually applied to various content production pipelines in society and the Internet, some potential risk concerns are also worthy of consideration. Some studies have found that AI-generated content has risks in many aspects, such as discrimination, privacy leakage, ethics, and safety (Deshpande et al., 2023; Zhuo et al., 2023; Chen and Shu, 2023; Jiang et al., 2023; Su et al., 2023). A recent study finds that texts generated by AI introduce bias in text retrieval, causing the retriever to rank them higher (Dai et al., 2023). This paper extends the study of this bias to cross-modal retrieval.
@@ 7. Conclusion
This paper delves into the exploration of the bias in text-image retrieval introduced by AI-generated images. We construct a reasonable benchmark to simulate the retrieval scenarios comprising both real images and AI-generated images. Experiments on this benchmark underscores that AI-generated images tend to be ranked higher by retrieval models, despite lacking more visually relevant semantics to the queries than real images. We define this bias as invisible relevance bias. This bias is prevalent across retrieval models with varying training data and architectures, including models trained from scratch and those pre-trained on massive image-text pairs, and encompassing dual-encoder and fusion-encoder models. Moreover, mixing AI-generated images into the training data makes the bias more serious, causing a vicious cycle where AI-generated images gain more exposure from massive data, increasing their likelihood of being mixed into retrieval model training, and exacerbating the bias further. To solve this, we propose an effective debiasing method to mitigate the invisible relevance bias. Then, we use our proposed debiasing method to reversely determine that the cause of invisible relevance is that the AI-generated images cause the image encoder to introduce additional information into their representation. This additional information can make the retriever estimate a higher relevance score. Findings in this paper reveal the potential impact of AI-generated images on text-image retrieval systems in the context of the rapid development of AIGC and have implications for further research.

### v4

@@ 3.4. More Serious Bias Caused by Training
Due to the wide distribution of AI-generated images on the internet and the bias in Section 3.3, AI-generated images are very likely to be mixed into the training data of retrieval models. This section delves deeper into the impact on retrieval performance and invisible relevance bias when AI-generated images are mixed into the training of retrieval models. The experimental results reveal a vicious cycle of falling into more serious invisible relevance bias. Specifically, the invisible relevance bias of text-image retrieval models causes the AI-generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of retrieval models, leading to more serious bias and forming a vicious cycle. This ultimately results in users’ search results being surrounded by AI-generated images.
Specifically, we explore the impact of the training mixed with generated images on retrieval by incorporating varying ratios of generated images into the training data. To ensure an accurate assessment, our experiments focus on the model trained from scratch (VSE). This is because pre-trained vision-language models have been pre-trained on massive real text-image pairs, which introduce additional biases in evaluating the impact of specific ratios of generated images in the training data. We reconstruct the training set of Flicker30k by replacing a certain ratio of real images with AI-generated images. For each real image ($I_{r}$) in the training set, we use the method in Section 2 to generate its corresponding AI-generated image ($I_{g}$). Then, we introduce the ratio $\alpha$, which means that in our reconstructed training data, the paired images for $\alpha$ percentage of captions are AI-generated images, and for $(100-\alpha)$ percentage of captions are real images. We change the ratio while keeping the total number of training samples unchanged. We evaluate the performance of the trained model on the test set of Flicker30k+AI (in-domain setting) and MSCOCO+AI (out-of-domain setting).
@@
Figure 2. Assessment results on the training set mixed with AI-generated images. We change the ratio of AI-generated images in the datasets (X-axis) while keeping the total number of training samples unchanged. The model is tested on the test set of Flicker30k+AI (in-domain) and MSCOCO+AI (out-of-domain) respectively that we constructed in Section 2.
@@
Figure 2 illustrates that as the ratio increases, the ranking disparity between generated images and real images widens, with the retrieval model exhibiting a greater inclination to rank generated images higher ($\rm{Relative}\triangle$ decreasing). Meanwhile, Figure 3 demonstrates that with an increasing ratio, the discrepancy in score distribution between generated images and real images increases, and the scores of AI-generated images gradually become greater and greater than those of real images. In both in-domain and out-of-domain settings, the following conclusions can be drawn from these results: (1) Retrieval models trained on the datasets mixed with AI-generated images exhibit more serious invisible relevance bias. (2) The invisible relevance bias tends to become more serious as the ratio of AI-generated images in the training data increases. (3) As the ratio of AI-generated images in the training data increases, the retrieval performance on real images is gradually damaged.
@@ 4. Our Method in debiasing
This section proposes an effective method to alleviate the bias.
@@ 4.1. Design of Debiasing Method
Our method alleviates the retrieval model’s preference bias for generated images by measuring and optimizing the additional relevance score of generated images in training. Given a caption-image pair $(C_{i},I^{r}_{i})$ in the training set, the training process for both dual-encoder and fusion-encoder text-image retrieval models can be formulated as estimating the relevance score $s$ between $C_{i}$ and $I^{r}_{i}$, and using contrastive loss or regression loss as the optimization objective to adjust $s$, which can be described as:
@@
in which $\mathbb{A}$ is the set of all samples in the training data. For the sample $i$, $s_{i}$ is the estimated score and $y_{i}$ is the label.
4.2. Evaluation of Debiasing
@@
Evaluation of our debiasing method focuses on: (1) How it affects the ranking difference between real and AI-generated images. (2) How it affects the distribution of caption-image relevance scores.
Table 5 shows how the $\rm{Relative\triangle}$ and retrieval performance changes with the sampling probability $\beta$. The results indicate that: (1) Our method not only effectively alleviates the retrieval model’s preference for AI-generated images, but also makes real images ranked significantly higher than AI-generated images. (2) As the sampling probability $\beta$ increases, real images are ranked higher and higher than AI-generated images. (3) When $\beta$ is $0.5$, retrieval models can achieve a fair ranking between real and generated images with little bias. (4) Our method improves the retrieval performance. It is because, in our method, AI-generated images can be seen as the hard negatives in training, which enhances the ability of the retrieval model to distinguish the images with very similar semantics.
Figure 4 shows the distribution of the caption-image relevance scores estimated by retrieval models with different sampling probability $\beta$. The results indicate that our method effectively reduces the relevance scores between captions and generated images estimated by the retrieval model. With the increase in sampling probability ($\beta$), the disparity in score distribution between generated images and real images expands, and the scores of real images gradually become greater and greater than those of generated images.
@@ 5. Causes of Invisible Relevance Bias
In this section, we use the debiased model proposed in Section 4 to reversely analyze the causes of the invisible relevance bias. Specifically, we extract the transformations performed by the debiased retriever on AI-generated images. The reverse process of these transformations can be seen as the cause of invisible relevance bias. We also design the experiments to support this point of view.
@@ 5.1. Transformations in Debiasing
To simplify the analysis process, we focus on the dual-encoder retrieval model that estimates the relevance score by computing vector similarity between text and image representations. For the same text, image representation can directly affect the estimation of the relevance, so we use image representation as the main object in our analysis, which can help us to find the most direct causes of invisible relevance bias. Specifically, we analyze the difference between image representations encoded by the original retriever and debiased retriever and extract the transformations of debiasing from this difference. Then, we reverse these transformations to explore the causes of the invisible relevance bias.
Given the image encoder $v(\cdot;\theta)$ of original retriever, the image encoder $v^{d}(\cdot;\theta^{\prime})$ of our debiased retriever, and an AI-generated image $I^{g}$, we can get the representations of $I^{g}$ encoded by $v(\cdot;\theta)$ and $v^{d}(\cdot;\theta^{\prime})$ respectively as $\bm{r}=v(I^{g};\theta),\bm{r^{d}}=v^{d}(I^{g};\theta^{\prime}).$ For the normalized text representation $\bm{t}=[t_{1},t_{2},...,t_{n}]$, image representation $\bm{r}=[r_{1},r_{2},...,r_{n}]$ from the original retriever and image representation $\bm{r^{d}}=[r^{d}_{1},r^{d}_{2},...,r^{d}_{n}]$ from the debiased retriever, the relevance $score(\bm{t},\bm{r})$ and $score(\bm{t},\bm{r^{d}})$ can be calculated as:
@@
$\bm{p}$ $=[\triangle r_{1},\triangle r_{2},...,\triangle r_{n}]=[p_{1},p_{2},...,p_{n}],$
@@ 5.2. Reversing the Transformations
We can reverse this process to get the causes of the invisible relevance bias. This is because the transformation made by a debiased retriever to make a biased AI-generated image become unbiased is exactly the difference between the biased AI-generated images and the real images. That is, the reason why an unbiased representation $\bm{r^{d}}$ becomes the representation $\bm{r}$ with bias is that the reverse transformation ($-p_{i}$) is done on $r^{d}_{i}$ ($r_{i}=r^{d}_{i}-p_{i}$). Therefore, we conclude that an unbiased representaion $\bm{r^{d}}$ becomes the representaion $\bm{r}$ with bias because $\bm{r^{d}}$ is added bitwisely by a vector $\bm{-p}=[-p_{1},-p_{2},-p_{3},...,-p_{n}]$. Combining the phenomenon shown in Figure 5 that there is consistency in the transformations vector $\bm{p}$ on AI-generated images with very different semantics, we can get the causes of invisible relevance bias in AI-generated images: AI-generated images cause the image encoder in the retriever to embed additional information to their representations. This additional information is the direct cause of invisible relevance bias and can be obtained by the difference in image representations between original and debiased retrievers. This information has the following three characteristics: (1) This information cannot be reflected in a visible way, but can only be embedded by neural network-based models. (2) When this information is embedded into the representation of the image, it can amplify the query-image relevance to produce a higher score. (3) This information has a certain consistency in AI-generated images with different semantics. This information may be like the watermark that is universal information for the image generation model and can be expressed by neural network visual models such as image encoders. We design the experiments to support these three points.
@@ 6.1. Artificial Intelligence Generated Content
With the development of generation models, AI-generated content (AIGC) becomes more realistic and less discernible (Agnese et al., 2020; Cao et al., 2023; Wu et al., 2023). The use of AI to generate content has greatly changed the way of content generation. It reduced the cost of content generation and improved the efficiency than traditional compared to traditional content generation pipeline biased on humans (Hanley and Durumeric, 2023; Spitale et al., 2023). For example, Stable Diffusion (Yang et al., 2023; Croitoru et al., 2023) and DALL-E-3 (Betker et al., 2023) can generate high-quality images by simply following a short text description of uses’ input. However, as AIGC is gradually applied to various content production pipelines in society and the internet, some potential risk concerns are also worthy of consideration. Some studies have found that AI-generated content has risks in many aspects, such as discrimination, privacy leakage, ethics, and safety (Deshpande et al., 2023; Zhuo et al., 2023; Chen and Shu, 2023; Jiang et al., 2023; Su et al., 2023). A recent study finds that texts generated by AI introduce bias in text retrieval, causing the retriever to rank them higher (Dai et al., 2023). This paper extends the study of this bias to cross-modal retrieval.
@@ 7. Conclusion
This paper explores the impact on text-image retrieval introduced by AI-generated images. We construct a reasonable benchmark to simulate the retrieval scenarios comprising both real images and AI-generated images. Experiments on this benchmark underscore that AI-generated images tend to be ranked higher by retrieval models, despite lacking more visually relevant semantics to the queries than real images. We define this bias as invisible relevance bias. This bias is prevalent across retrieval models with varying training data and architectures. Moreover, mixing AI-generated images into the training data makes the bias more serious, causing a vicious cycle where AI-generated images gain more exposure from massive data, increasing their likelihood of being mixed into retrieval model training, and exacerbating the bias further. To solve this, we propose an effective debiasing method to mitigate the invisible relevance bias. Then, we use our proposed debiasing method to reversely determine that the cause of invisible relevance is that the AI-generated images cause the image encoder to introduce additional information into their representation. This additional information can make the retriever estimate a higher relevance score. Findings in this paper reveal the potential impact of AI-generated images on text-image retrieval systems in the context of the rapid development of AIGC and have implications for further research.
Acknowledgements.
This work was supported by the National Key R&D Program of China (2022YFB3103700, 2022YFB3103704), the National Natural Science Foundation of China (NSFC) under Grants No. 62276248 and U21B2046, and the Youth Innovation Promotion Association CAS under Grants No. 2023111.